# Feynman integrals at large loop order and the $\log$-$\Gamma$ distribution

### v2

@@ 3.2 Estimating moments and the primitive $\phi^{4}$ beta function
In our data (see Table 1), we observed that $\lambda>3$ for $L\leq 17$. Hence, up to this loop order, we may assume that $\langle\mathcal{P}(G)^{2}\rangle_{L}$ is finite and that we can safely estimate $\langle\mathcal{P}(G)\rangle_{L}$ using the central limit theorem by computing the average of all computed values of $\mathcal{P}(G)$. The central limit theorem has the advantage over Eq. (8) of providing accurate results if the data does not exactly follow distribution (2). The uncertainty of this estimate is computed as usual by dividing the sample variance by the number of samples and taking the square root. The results of the estimation are listed in Table 1. We also estimated $\langle{\mathcal{P}}(G)\rangle_{L}$ using Eq. (8). The discrepancy between both ways to estimate this expectation value decreases with the loop order $L$. At 17 loops, both methods give confidence intervals that overlap almost perfectly.
@@
Via (7), we can translate our estimates for $\langle\mathcal{P}(G)\rangle_{L}$ into the primitive contribution to the $\phi^{4}$ beta function. The necessary values of $Z_{L}$ were calculated using renormalized 0-dimensional QFT technology [47, §6.3] or [48] (see also [49] for a deeper analysis of similar normalization factors from 0-dimensional quantum field theory). Table 1 also includes the resulting estimates for $\beta_{L+1}^{\rm prim}$. These agree with the data for $L\leq 11$ in [23, Table XIII] and confirm estimates in [31, Table 14] up to 17 of 18 loop orders.
@@ 3.3 Extrapolating $\beta_{L}^{\rm prim}$ to all loop orders using instanton input
Unfortunately, our data for $\beta^{\mathrm{prim}}_{L}$ does not provide compelling evidence for the asymptotic behavior (9). We find, via goodness-of-fit estimates, that our data for $\beta^{\mathrm{prim}}_{L}$ is compatible with the following Ansatz, which Eq. (9) directly inspires:
@@
By varying the fit range and the cut-off point of the expansion in powers of $1/L$ in (10), we obtained various estimates for the value of $c_{0}$. With reasonable choices for both the fit range and the cut-off point, $c_{0}$ consistently falls into the range of $0.055\pm 0.015$. Even though the predicted value of $C\approx 0.024$ lies outside this range, the large margin of error does not allow for a clear verdict on the validity of (9). Moreover, the results of [49] show that asymptotic estimates such as ours, which rely on low-order computations, suffer from various biases. The authors highlight a particular case relevant to $\phi^{4}$ theory, where the true asymptotic behavior only becomes apparent if the first 25 perturbative coefficients are known.
@@ 4 Conclusion
The limiting distribution has three remaining parameters, $\alpha,\lambda,$ and $\mathcal{P}_{0}$, that we fix by fitting at each loop order. It would be highly beneficial to find explicit limiting laws for these three parameters, i.e. to find the asymptotic behavior of these numbers for $L\rightarrow\infty$.
@@
We used our data to compute the primitive contribution to the $\phi^{4}$ beta function up to 17 loops confirming previous results by Balduf [31]. We discussed the extrapolation of this contribution to all loop orders. It is conjectured that the primitive beta function equals the $\phi^{4}$ beta function in the minimal subtraction scheme when $L\rightarrow\infty$, so our extrapolation provides a (conjectured) estimate of also this beta function at infinite loop order. This second beta function has numerous phenomenological applications via the Wilson–Fischer approach to critical phenomena. We postpone the discussions of the phenomenological implications of our findings (e.g., on the critical exponents of the $D=3$ Ising model) to future work.
Unfortunately, reaching a verdict of full or only partial agreement between our data and predictions from instanton computations is still impossible. Our limited data, the resulting poor fit quality, and the large number of sources for numerical perturbations do not allow a clear conclusion. More data at an even higher loop order seems necessary to complete the picture. The key limiting factor of our computations was the memory requirements of the tropical sampling implementation [45]. Harnessing more properties of the Hepp bound from [30] and the (tropical) geometry of Feynman integrals might reduce these requirements and make higher loop orders accessible (see [26, §8.1 and §8.3]).

### v3

@@ 3.2 Estimating moments and the primitive $\phi^{4}$ beta function
In our data (see Table 2), we observed that $\lambda>3$ for $L\leq 17$. Hence, up to this loop order, we may assume that $\langle\mathcal{P}(G)^{2}\rangle_{L}$ is finite and that we can safely estimate $\langle\mathcal{P}(G)\rangle_{L}$ using the central limit theorem by computing the average of all computed values of $\mathcal{P}(G)$. The central limit theorem has the advantage over Eq. (9) of providing accurate results even when, at a given finite loop order, the empirical distribution deviates slightly from (2). The uncertainty of this estimate is computed as usual by dividing the sample variance by the number of samples and taking the square root. The results of the estimation are listed in Table 2 under the column $\langle{\mathcal{P}}(G)\rangle_{L}^{\rm CLT}$. We also estimated $\langle{\mathcal{P}}(G)\rangle_{L}$ using Eq. (9), and the results of the estimation are listed in Table 2 under the column $\langle{\mathcal{P}}(G)\rangle_{L}^{\log\text{-}\Gamma}$. The discrepancy between both ways to estimate this expectation value decreases with the loop order $L$. At 17 loops, both methods give confidence intervals that overlap well.
@@
Via (7), we can translate our estimates for $\langle\mathcal{P}(G)\rangle_{L}$ into the primitive contribution to the $\phi^{4}$ beta function. The necessary values of $Z_{L}$ were calculated using renormalized 0-dimensional QFT technology [52, §6.3] or [53] (see also [54] for a deeper analysis of similar normalization factors from 0-dimensional quantum field theory). Table 2 also includes the resulting estimates based on the values $\langle{\mathcal{P}}(G)\rangle_{L}^{\rm CLT}$ for $\beta_{L+1}^{\rm prim}$. These agree with the data for $L\leq 11$ in [23, Table XIII] and confirm estimates in [33, Table 14] and [54, Table 4] up to 17 of 18 loop orders.
Extending the results of [33] to the context of Feynman periods sampled proportionally to their symmetry factors, we also empirically investigated the central moments of the distribution of $\mathcal{P}(G)$ using our dataset. For $k\geq 2$, the central moments $\mu_{k}$ are defined by
$\mu_{k}=\langle\left(\mathcal{P}(G)-\langle\mathcal{P}(G)\rangle_{L}\right)^{k}\rangle_{L}.$ (10)
As discussed above, applying the central limit theorem to estimate our quantity of interest, $\langle\mathcal{P}(G)\rangle_{L}$, requires that the second moment $\mu_{2}$ is finite. Further, to estimate the accuracy of this estimate itself, the variance of this second moment, $\mathrm{Var}[\mu_{2}]$, needs to be finite as well. Standard arguments in statistics (see, e.g., [55, Sec. 6.10]) show that finiteness of this variance is implied by the finiteness of the third and fourth moments $\mu_{3}$ and $\mu_{4}$. Hence, we computed the moments $\mu_{2},\mu_{3},$ and $\mu_{4}$ from our data. The results are presented in Table 3 and Figure 4. Up to a different normalization convention, our results are consistent with those reported in [33, Table 7].
Following the approach of [33], we fit the computed moments using the function family:
$f(L;a,b)=e^{a}L^{b},$ (11)
where $a$ and $b$ are fit parameters. The results of these fits are listed in Table 4 and illustrated as colored curves in Figure 4. These fits show that the moments $\mu_{k}$ rapidly increase with the loop order $L$. Qualitatively, our fit results are also consistent with the behavior reported in [33] for uniformly sampled graphs.
The rapid growth of these moments suggests that estimating $\langle\mathcal{P}(G)\rangle_{L}$ with the central limit theorem will become more and more difficult for larger $L$. Estimating this average using the $\log$-$\Gamma$ hypothesis by fitting the parameters $\alpha,\lambda$ and $\mathcal{P}_{0}$ and using formula (9) on the other hand can be expected to stay feasible at arbitrary large loop orders. The turning point, where estimating the average using the $\log$-$\Gamma$ hypothesis gives as accurate results as estimating via the central limit theorem, seems to be at $L\approx 17$ as indicated by the data in Table 2.
$L$ $\mu_{2}$ $\mu_{3}$ $\mu_{4}$
$10$ $2.4121\cdot 10^{6}$ $4.6142\cdot 10^{9}$ $2.9535\cdot 10^{13}$
$11$ $1.4479\cdot 10^{7}$ $8.7965\cdot 10^{10}$ $1.3856\cdot 10^{15}$
$12$ $7.9490\cdot 10^{7}$ $1.4266\cdot 10^{12}$ $5.6400\cdot 10^{16}$
$13$ $3.9723\cdot 10^{8}$ $1.9882\cdot 10^{13}$ $2.0226\cdot 10^{18}$
$14$ $1.8339\cdot 10^{9}$ $2.3361\cdot 10^{14}$ $5.7832\cdot 10^{19}$
$15$ $7.9001\cdot 10^{9}$ $2.5503\cdot 10^{15}$ $1.5655\cdot 10^{21}$
$16$ $3.1370\cdot 10^{10}$ $2.3915\cdot 10^{16}$ $3.5246\cdot 10^{22}$
$17$ $1.2168\cdot 10^{11}$ $2.0207\cdot 10^{17}$ $6.1323\cdot 10^{23}$
Table 3: Empirically measured moments $\mu_{k}$ at different loop orders $L$.
Figure 4: Empirical moments $\mu_{k}$ of the distribution of $\mathcal{P}(G)$ as a function of the loop order $L$. The curves are fits to the data points using the function family (11).
$a$ $b$
$\mu_{2}$ 32.47(57)-32.47(57) $20.42(22)$
$\mu_{3}$ 54.44(78)-54.44(78) $33.22(30)$
$\mu_{4}$ 71.9(1.6)-71.9(1.6) $44.54(63)$
Table 4: Best-fit parameters $a$ and $b$ for the empirical moments $\mu_{k}$ using the function family (11).
@@ 3.3 Extrapolating $\beta_{L}^{\rm prim}$ to all loop orders using instanton input
Unfortunately, our data for $\beta^{\mathrm{prim}}_{L}$ does not provide compelling evidence for the asymptotic behavior (12). The main obstacle towards a clear verdict seems not to be the limited accuracy, but the still relatively low loop order of our data points. We find, via goodness-of-fit estimates, that our data for $\beta^{\mathrm{prim}}_{L}$ is compatible with the following Ansatz, which Eq. (12) directly inspires:
@@
By varying the fit range and the cut-off point of the expansion in powers of $1/L$ in (13), we obtained various estimates for the value of $c_{0}$. With reasonable choices for both the fit range and the cut-off point, $c_{0}$ consistently falls into the range of $0.055\pm 0.015$. Even though the predicted value of $C\approx 0.024$ lies outside this range, the large margin of error does not allow for a clear verdict on the validity of (12). Moreover, results of Balduf and Thürigen [54] show that asymptotic estimates such as ours, which rely on low-order computations, suffer from various biases. The authors highlight a particular case relevant to $\phi^{4}$ theory, where the true asymptotic behavior only becomes apparent if the first 25 perturbative coefficients are known.
@@ 4 Conclusion
The limiting distribution has three remaining parameters, $\alpha,\lambda,$ and $\mathcal{P}_{0}$, that we fix by fitting at each loop order. It would be highly beneficial to find explicit limiting laws for these three parameters, i.e. to find the asymptotic behavior of these numbers for $L\rightarrow\infty$. Here, we only studied the behaviour of these parameters qualitatively by making shifted power-law fits.
@@
We used our data to compute the primitive contribution to the $\phi^{4}$ beta function up to 17 loops confirming previous results by Balduf [33]. We discussed the extrapolation of this contribution to all loop orders. It is conjectured that the primitive beta function equals the $\phi^{4}$ beta function in the minimal subtraction scheme when $L\rightarrow\infty$, so our extrapolation provides a (conjectured) estimate of also this beta function at infinite loop order. The beta function in the minimal subtraction scheme has numerous phenomenological applications via the Wilson–Fischer approach to critical phenomena. We postpone the discussions of the phenomenological implications of our findings (e.g., on the critical exponents of the $D=3$ Ising model) to future work.
Unfortunately, reaching a verdict of full or only partial agreement between our data and predictions from instanton computations is still impossible. Our limited data, the resulting poor fit quality, and the large number of sources for numerical perturbations do not allow a clear conclusion. More data at an even higher loop order seems necessary to complete the picture. The key limiting factor of our computations, which kept us from studying even higher loop orders, was the memory requirements of the tropical sampling implementation [50]. Harnessing more properties of the Hepp bound from [30] and the (tropical) geometry of Feynman integrals might reduce these requirements and make higher loop orders accessible (see [26, §8.1 and §8.3]).